# A Distributed ADMM-based Deep Learning Approach for Thermal Control in Multi-Zone Buildings

### v1

@@ II Problem formulation
The sharing problem can be solved using Alg. 1. We distinguish two main sub-problems in this algorithm. The first one is the local controller problem (line 5) and the second one is the coordinator problem (line 7).
As stated in the following theorem, under some Lipschitz regularity assumptions and the right choice of $\rho$, this algorithm is guaranteed to converge although the problem is not convex. We will assess the quality of the solutions experimentally in section V.
@@
The Alg. 1 converges to the set of stationary solutions of problem II :
@@
Evaluating the models of each dynamics $q_{i}$ is computationally expensive. To leverage the benefits of the decomposition, the models’ evaluation should be placed in the for loop and computed in parallel. In Alg. 1 the models’ evaluation is in the $\ell$ function, in the coordinator problem.
We can explicit the relationship between the setpoints and the related power consumption (4) as a constraint in the LC problem.
{mini!}
[…]
and consider the surrogate coordinator problem:
{mini!}
[3] ¯u~ℓ(¯u) + ∑_i=1^N λ_i^T (¯u_i - u_i) + ρ—— ¯u_i - u_i ——_2^2
where ~$\tilde{\ell}(\bar{\textbf{u}})=\left\lVert A\bar{\textbf{u}}-\bar{P}^{tot}\right\rVert$. This formulation may be derived from the same decomposition steps as the ones above, but using the following equation rather than eq. (4):
@@
Eq. (6) bounds the state and power usage to a setpoint. The problem is that $\psi$ is not an injective application. For instance, if the temperature in a zone is $20^{\circ}C$, temperature setpoints below $20^{\circ}C$ will all yield a heating power of $0W$. Such an application is impossible to learn accurately in practice. That said, depending on the type of setpoint tracking system, how the power is measured and the operation regime, one may find conditions for $\psi$ to satisfy the requirements of Theorem 1. For instance, a proportional controller that operates with a zone’s temperature around the setpoint or an ON/OFF controller used with Pulse Width Modulation would give good properties to $\psi$.
Instead of enumerating such conditions, that are hard to verify in practice, we choose to provide a theoretical analysis for the problem formulation (4). This formulation is more natural and indicates the form of the ADMM decomposition for the surrogate problem. As shown in Section V, the surrogate formulation yields efficient computation and shows excellent convergence properties experimentally. In addition, it provides a good intuition about the underlying mechanism: the coordinator (i.e. problem (II.1)) is choosing the amount of power for each room and each timestep, to ensure that the power constraint is always satisfied. The LCs (i.e. problems (II.1)) must then find the temperature setpoints to track this power target while preserving the users’ comfort. Even if multiple setpoints may satisfy the constraints II.1 and II.1, the lowest setpoint changes will be chosen because of the term $g(\Delta_{i})$ is the objective function.
The control architecture is summarized in Fig. 1.
Figure 1: Two layer hierarchical control architecture. The LCs observe the state of their zone and plan for the best setpoints while communicating with the Coordinator to enforce a maximum constraint on power usage. Setpoints are sent to the existing setpoint trackers that act on each zone.
III Local Controllers and Zone Modelling
@@
To successfully plan for the best setpoint changes without implementing any actions, LCs must be able to predict the evolution of the system. In this section, we present two methods for the modeling of a thermal zone: Recurrent State Space Models (RSSM) and State Space Models (SSM).
III-A Reccurent State Space Model
@@
    A representation model (encoder), which maps the observations $o_{t}$ to the latent state $s_{t}$ of the system. The use of an encoder is important because the system is partially observed. One can include lags of power and temperature in the observation to build a better representation of the state and cope with partial observability.
@@
    A transition model, that predicts the evolution of the latent space. Note that this transition model is composed of two parts: (1) a deterministic state model which predicts the evolution of the deterministic part of the state $h_{t}=f(h_{t-1},s_{t-1},\delta_{t-1})$ and (2) a stochastic state model which samples the next latent state according to $h_{t}$ : $s_{t}\sim p(s_{t}|h_{t})$. The latter part is important in our case as it accounts for the perturbations due to the occupancy and neighboring zones.
@@
All these components are represented by neural networks. We use the same set of parameters $\theta$ to learn the observation model, the transition model, and the representation model. The objective is to maximize the variational bound as proposed in [32]. The transition model allows to predict one timestep ahead in the latent space. Multiple steps prediction is achieved by applying the model recursively. This is convenient as the user may want to change the prediction horizon without having to retrain the model. The latent space embeds the inner state of the room, mainly composed of the power usage and the room’s temperature. This state is influenced by exogenous perturbations such as weather and the temperature setpoint. To not overload the transition model, we do not include these disturbances in the latent space as it is usually done with this type of model. Instead, we feed the transition model with the previous latent space as well as forecasts of the disturbances, as pictured in Fig. 3. The advantages of this approach are that a single weather prediction system is required for the entire building. Since the RSSM of each zone does not have to predict the complex evolution of the weather on top of the evolution of the room’s state, it can provide better predictions using a smaller network model. In addition, some disturbances such as temperature setpoint schedules may be known beforehand and can be passed exactly to the transition model at each prediction step.
@@
III-B State Space Models
The State Space Model (SSM) has the same structure as the RSSM, but it does not include the stochastic part. The transition model is a GRU cell that takes as input the disturbances (i.e. the weather and the setpoints). The latent state $s_{t}$ is carried by the hidden state of the GRU cell. The model is trained to minimize the mean squared error between the predicted and actual powers.
IV Distributed Planning Networks
In this section, we present how the models of each LC are combined with the coordinator using ADMM. The proposed algorithm is called Distributed Planning Networks as it embeds the idea of the Planning Networks in a distributed architecture. Fig. 2 and Alg. 3 summarize the approach.
The coordinator problem is a convex penalized least square problem. It can be solved easily with existing solvers. The challenge is more on the side of the LC problem (II.1).
To efficiently find a solution to the non-convex LC problem, we propose two approaches. The first one is a random search for the best temperature setpoints. The second one is a shooting method, where we differentiate through the prediction model and use the gradient to guide the search. Once the best setpoint changes have been found for the upcoming horizon, only the setpoint changes of the first timestep are implemented. The process is then repeated to allow for re-planning and correction of the control.
IV-A Planning by Random Search with RSSM
The random search is conducted as follows:
@@ IV-A Planning by Random Search with RSSM
    Sample the trajectories corresponding to each $\boldsymbol{\Delta}_{i}$ using the RSSM. For each $\boldsymbol{\Delta}_{i}$, the RSSM outputs a distribution for the trajectory of states. For one choice of $\boldsymbol{\Delta}_{i}$, multiple states trajectories may be sampled from the distribution to account for uncertainties.
@@
    Select $\boldsymbol{\Delta}_{i}$ providing the lowest score as defined by (II.1).
@@
While using ADMM, the steps presented above need to be repeated at each iteration. The second step is computationally expensive as it requires many forward passes in the transition and observation networks. However, only the third step depends on the ADMM iteration. To limit the computational time, the trajectories (i.e. the two first steps) are pre-computed and stored before the start of ADMM. Note that these two steps can be performed in parallel for each zone. At each ADMM iteration, the score of each trajectory is then computed from the pre-computed trajectory with respect to the current power targets $\bar{\textbf{u}}_{i}$.
IV-B Planing by Shooting with SSM
The SSM is fully differentiable. One can formulate the problem (II.1) as an Initial Value Problem and use single shooting to minimize the running cost given by the objective function. The equality constraints representing the dynamics are implicitly satisfied by the prediction model. To handle the inequality constraint (II.1), we use projected gradient descent to keep the setpoint changes within the desired interval. This approach assumes continuous values for the setpoints. In practice, we round the values at the resolution of the setpoint tracker of the zone afterwards. A stopping criterion on the number of decimals changing in $\Delta_{i}$ at each iteration is used to make the approach computationally efficient.
V Results
To test the proposed approach, we use the low-rise apartment building, one of the Department Of Energy (DOE) residential reference buildings. The building is modeled in EnergyPlus, a software that provides thermal simulations for buildings with state-of-the-art accuracy. The building is composed of three floors of 6 apartments each. Apartments are equipped with independent HVAC systems, temperature setpoints schedules, occupancy schedules, and appliance usage schedules. Simulations are run using Typical Meteorological Year (TMY) files coming from the DOE climate zone 7A, representing the weather in the North-East of the United States of America and South-East of Canada. In each apartment, the user setpoint $T_{i}^{sp}$ is kept constant, but may vary between $19^{\circ}C$ and $22^{\circ}C$ from one apartment to the other. For the purpose of this study, we are considering heating periods, and hence we modify the default HVAC systems to electric heating-only systems with an efficiency of one. The EnergyPlus timestep was set to 15 minutes. The prediction horizon is set to one or two hours and the ADMM timestep is 30 minutes, meaning we can re-plan to change the setpoints every 30 minutes.
V-A Models accuracy
zone RSSM SSM zone RSSM SSM
1h ahead forecast 1 185 78 10 125 89
2 129 79 11 158 140
3 181 88 12 127 71
4 136 67 13 177 92
5 122 55 14 134 94
6 130 68 15 164 105
7 137 94 16 152 69
8 144 116 17 128 57
9 159 107 18 134 84
zone RSSM SSM zone RSSM SSM
2h ahead forecast 1 184 80 10 98 82
2 120 71 11 144 133
3 182 88 12 109 72
4 134 63 13 173 103
5 104 51 14 112 96
6 127 66 15 159 110
7 127 105 16 122 71
8 131 121 17 105 79
9 152 108 18 109 102
TABLE I: Models scores (MAE, Watt) for one hour and two hours ahead predictions during a three weeks test period. For the entire building’s power forecast, the RSSM has a MAPE of $5.81\%$ for one hour ahead and $6.20\%$ for two hours ahead. The SSM has a MAPE of $3.72\%$ and $4.58\%$ respectively.
[…]
First, we test the accuracy of the prediction models. Both RSSM and SSM are trained offline for each zone of the building. The training data are generated by running simulations with random temperature setpoints changed every period of random duration, ranging from 1h to 24h. The observation of a zone contains the zone’s current temperature as well as the power consumption of each timestep for the last hour. The output of the representation model is the power consumption of the next timestep.
At each timestep, the zone’s observation is completed with disturbances inputs. Disturbances include the temperature setpoint and the temperature setpoint changes that are known beforehand. As one may observe in Fig. 4, the changes in the temperature setpoints cause steep variations in the power. To accurately predict these changes, a key feature added is the difference between setpoint changes from one timestep to the next i.e. $\delta_{t,i}-\delta_{t-1,i}$. The disturbances also include the weather, composed of the outdoor temperature and solar radiation. For simplicity, we did not build a weather forecast model. Instead, we used perfect weather information and added random noise to account for prediction errors.
Table I shows the Mean Absolute Error (MAE) of each zone for one-hour and two-hours ahead forecasts. The SSM performs consistently better than the RSSM, but both methods give accurate predictions. For the entire building’s power forecast, the Mean Average Percentage Errors (MAPE) of the SSM are $3.72\%$ and $4.58\%$ for one-hour and two-hours ahead predictions. The MAPE of the RSSM are $5.81\%$ and $6.20\%$ respectively. These predictions were made without observing the states of the neighboring zones, as stated in the assumption of the model used in (II).
V-B Distributed Planning Networks and Power Control
To test the Distributed Planning Networks algorithm, we propose to implement it for peak load reduction. We consider a Demand Response (DR) program where the utility sends an early notification to the building’s operator to reduce the power demand during peak demand events. We assume that the events occur between $5AM$ and $9AM$ every morning over a period of three weeks. We also assume that this reduction implies a limit on the heating demand that should not exceed $\textbf{P}^{max}_{t}=40kW$ for all timesteps $t$ during DR events. For each zone, we set $-2^{\circ}C\leq\boldsymbol{\Delta}_{t}\leq 0^{\circ}C$, with an action step of $0.25^{\circ}C$.
2:  Prediction models $\text{Prediction models }\phi_{1},\ldots,\phi_{N}$
3:  Parameters : $\text{Parameters : }\nu,\boldsymbol{\Delta}_{min},\boldsymbol{\Delta}_{max},H$
4:  Initialize $\textbf{P}^{tot}\in\mathbb{R}^{H}$
5:  for i=1, …, N do
6:     $P_{normal,i}=\phi_{i}(Q(s_{t},\boldsymbol{\Delta}_{max}))$
7:     $P_{min,i}=\phi_{i}(Q(s_{t},\boldsymbol{\Delta}_{min}))$
8:  end for
9:  $\textbf{P}_{normal}=\sum_{i=1}^{N}P_{normal,i},\text{ }\textbf{P}_{min}=\sum_{i=1}^{N}P_{min,i}$
10:  if $\textbf{P}_{normal}\leq(1-\nu)\textbf{P}^{max}$ then
11:     $\delta_{t,i}=0\text{ }\forall t,i$
12:     DO ADMM = FALSE
13:  else if $\textbf{P}_{min}>(1-\nu)\textbf{P}^{max}$ then
14:     $\delta_{t,i}=-2\text{ }\forall t,i$
15:     DO ADMM = FALSE
16:  else
17:     $\textbf{P}^{tot}=clip(\textbf{P}_{normal},(1-\nu)\textbf{P}^{max})$
18:     DO ADMM = TRUE
19:  end if

### v2

@@ II-B Distributed Sharing Algorithm
The sharing problem can be solved using Algorithm 1. We distinguish two main sub-problems in this algorithm. The first one is the Local Controller (LC) problem (line 5) and the second one is the coordinator problem (line 7).
As stated in the following theorem, under some Lipschitz regularity assumptions and the right choice of $\rho$, this algorithm is guaranteed to converge although the problem is not convex. We will assess the quality of the solutions experimentally in section IV. A proof of the theorem is given the Appendix A-A.
@@
The Algorithm 1 converges to the set of stationary solutions of Problem (14):
@@
Querying the zones’ dynamics $q_{i}$ is computationally expensive. To leverage the benefits of the decomposition, the dynamics’ evaluation should be placed in the for loop and computed in parallel. In Algorithm 1 the dynamics are called in the $\ell$ function, in the coordinator problem. It means that the local controllers are sending their temperature setpoint schedules to the coordinator which has to query all the zones’ dynamics to compute the associated HVAC consumption. Algorithm 1 formulation is convenient for a theoretical analysis of the problem. In practice, however, it is more efficient to compute the HVAC consumption at the LC level and exchange power consumption messages with the coordinator. We thus explicit the relationship between the setpoints and the related power consumption given by Eq. (11) as a constraint in the LC problem and derive a surrogate LC problem:
$\underset{\boldsymbol{\Delta}_{i}}{\text{minimize}}\quad$ $g(\boldsymbol{\Delta}_{i})+\boldsymbol{\lambda}_{i}^{T}(\bar{\textbf{u}}_{i}-\textbf{u}_{i})+\rho||\bar{\textbf{u}}_{i}-\textbf{u}_{i}||_{2}^{2}$ (18)
subject to $u_{t,i}=\phi(Q_{t,i}(\boldsymbol{s}_{0,i},\boldsymbol{\Delta}_{i}))$ (19)
$\bar{u}_{t,i}=\phi(Q_{t,i}(\boldsymbol{s}_{0,i},\bar{\boldsymbol{\Delta}}_{i}))$ (20)
$m_{i}\leq\delta_{t,i}\leq M_{i}$ (21)
as well as a surrogate coordinator problem:
minimizeu¯~($\underset{\bar{\textbf{u}}}{\text{minimize}}\quad\tilde{\ell}(\bar{\textbf{u}})+\sum_{i=1}^{N}\boldsymbol{\lambda}_{i}^{T}(\bar{\textbf{u}}_{i}-\textbf{u}_{i})+\rho||\bar{\textbf{u}}_{i}-\textbf{u}_{i}||_{2}^{2}$ (22)
where ~$\tilde{\ell}(\bar{\textbf{u}})=\left\lVert A\bar{\textbf{u}}-P^{tot}\right\rVert$. In our implementation, these surrogate problems are used in place on the LC and coordinator problem in Algorithm 1. The resulting control architecture is summarized in Fig. 1.
Note that the surrogate formulation may be derived from the same decomposition steps as the ones presented in Section II-A, but using the following equation rather than Eq. (11):
@@
Eq. (23) bounds the state and power usage to a setpoint. The problem is that $\psi$ is not an injective application. For instance, if the temperature in a zone is $20^{\circ}C$, temperature setpoints below $20^{\circ}C$ will all yield a heating power of $0W$. Such an application is impossible to learn accurately in practice. That said, depending on the type of setpoint tracking system, how the power is measured and the operation regime, one may find conditions for $\psi$ to satisfy the requirements of Theorem 1. For instance, a proportional controller that operates with a zone’s temperature around the setpoint or an ON/OFF controller used with Pulse Width Modulation would give good properties to $\psi$. Instead of enumerating such conditions, that are hard to verify in practice, we provide a theoretical analysis for the Problem (12). This formulation arises naturally from Problem (1) and indicates the form of the ADMM decomposition for the surrogate problem. As shown in Section IV, the surrogate formulation yields efficient computation and shows excellent convergence properties experimentally. In addition, it provides a good intuition about the underlying mechanism: the coordinator (i.e. Problem (22)) is choosing the amount of power for each room and each timestep, to ensure that the power constraint is always satisfied. The LCs (i.e. Problems (18)) must then find the temperature setpoints to track this power target while preserving the users’ comfort. Even if multiple setpoints may satisfy the constraints Eq. (19) and (20), the lowest setpoint changes will be chosen because of the term $g(\Delta_{i})$ is the objective function.
Figure 1: Two layers hierarchical control architecture. The LCs observe the state of their zone and plan for the best setpoints while communicating with the coordinator to enforce a maximum constraint on power usage. Setpoints are sent to the existing setpoint trackers that act on each zone.
III Distributed Planning Networks
In this section, we present the methods used to solve the optimization problems of Algorithm 1. In practice, the dynamics $q_{i}$ of the zones are unknown. In this study, we consider deep learning models to learn the dynamics from data. Both a deterministic and a stochastic approach are used, as described in the first part of this section. The models are used for planning to solve the LC problem (18), yielding the Distributed Planning Networks (DPN) algorithm. In the second part this section, we present a deterministic and a stochastic version of the DPN algorithm.
III-A Deep Learning Models for Thermal Zones
@@
To plan for the optimal setpoint changes, LCs must be able to predict the evolution of the thermal zones’ state. In this section, we present two types of models for the zones: State Space Models (SSM) and Recurrent State Space Models (RSSM). These models use inputs that include the observation of the zone $\textbf{o}_{t,i}$, actions $\textbf{a}_{t,i}$ and disturbances $\textbf{d}_{t}$ like weather and calendar information, to predict the next state in the observable space. To not overload the prediction models, the disturbances are not predicted by the models but given as inputs, as pictured in Fig. 3. One advantage of this approach is that a single weather prediction system is required for the entire building, and the zones’ model can focus on predicting only the zones’ states. The zones are partially observable environments and making predictions in a latent space helps to cope with partial observability.
III-A1 State Space Models
The SSM has an encoder-decoder structure and the transition function is a recurrent cell that takes as input the actions and disturbances. The latent state $s_{t}$ is carried by the hidden state of the recurrent cell:
$\mathbb{s}_{t+1}$ $=f_{\boldsymbol{\theta}}(\mathbb{s}_{t},\mathbb{a}_{t},\mathbb{d}_{t})$ (24)
$\mathbb{o}_{t+1}$ $=\text{dec}_{\boldsymbol{\theta}}(\mathbb{s}_{t+1})$ (25)
$\mathbb{s}_{0}$ $=\text{enc}_{\boldsymbol{\theta}}(\mathbb{o}_{-1},\ldots\mathbb{o}_{-n})$ (26)
where enc is an recurrent neural network encoding lags of observations, dec is a neural network decoding the latent state and $f$ is a recurrent cell used as the transition function. Prior works such as [23] have used similar models for representing buildings’ dynamics. The model is trained to minimize the mean squared error between the predicted and actual powers on the entire prediction horizon.
III-A2 Reccurent State Space Model
@@
    A representation model (encoder), which maps the observations $\mathbb{o}_{t}$ to the latent state $\mathbb{s}_{t}$ of the system.
@@
    A transition model, that predicts the evolution of the latent space. Note that this transition model is composed of two parts: (1) a deterministic state model which predicts the evolution of the deterministic part of the state $\boldsymbol{h}_{t}=f(\mathbb{h}_{t-1},\mathbb{s}_{t-1},\delta_{t},\mathbb{d}_{t})$ and (2) a stochastic state model which samples the next latent state $\mathbb{s}_{t}\sim p(\mathbb{s}_{t}|\mathbb{h}_{t})$. The latter part is important in our case as it accounts for the perturbations due to the occupancy and neighboring zones.
@@
All these components are represented by neural networks. We use the same set of parameters $\theta$ to learn the observation model, the transition model, and the representation model. The objective is to maximize the variational bound as proposed in [32]. The transition model predicts one timestep ahead in the latent space. Multiple-step prediction is achieved by applying the model recursively.
@@
III-B Distributed Planning Networks Algorithms
In this section, we present practical methods for solving the LC and coordinator optimization problems introduced in Algorithm 1. The coordinator problem is a convex penalized least square problem. It can be solved easily with existing solvers. The challenge is more on the side of the LC problem (18). For this problem, the zones’ models are used to plan for the best setpoint. We propose two approaches: the first one is a shooting method, where we differentiate through the prediction model and use the gradient to guide the search; the second one is a random search for the best temperature setpoints. Once the best setpoint changes have been found for the upcoming horizon, only the setpoint changes of the first timestep are implemented. The process is then repeated to allow for re-planning and correction of the control.
The proposed algorithm is called Distributed Planning Networks (DPN) as it embeds Planning Networks in a distributed architecture. Fig. 2 and Algorithm 2 summarize the approach.
III-B1 Deterministic DPN (DDPN)
The SSM is fully differentiable. One can formulate the Problem (18) as an initial value problem and use single shooting to minimize the running cost given by the objective function. The equality constraints representing the dynamics are implicitly satisfied by the prediction model. To handle the inequality constraint (21), we use projected gradient descent to keep the setpoint changes within the desired interval. This approach assumes continuous values for the setpoints. In practice, we round the values at the resolution of the setpoint tracker of the zone afterward. A stopping criterion on the number of decimals changing in $\Delta_{i}$ at each iteration is used to make the approach computationally efficient.
III-B2 Stochastic DPN (SDPN)
The random search is conducted as follow:
@@ III-B2 Stochastic DPN (SDPN)
    Sample the trajectories corresponding to each $\boldsymbol{\Delta}_{i}$ using the RSSM. For each $\boldsymbol{\Delta}_{i}$, the RSSM outputs a distribution for the trajectory of states. For one choice of $\boldsymbol{\Delta}_{i}$, multiple state trajectories are sampled from the distribution to account for uncertainties. The trajectory with the maximum power consumption is kept to optimize on the worst-case scenario.
@@
    Select $\boldsymbol{\Delta}_{i}$ providing the lowest score as defined by (18).
@@
While using ADMM, the steps presented above need to be repeated at each iteration. The second step is computationally expensive as it requires many forward passes in the transition and observation networks. However, only the third step depends on the ADMM iteration. To limit the computational time, the trajectories (i.e. the two first steps) are pre-computed and stored before the start of the ADMM iterations. Note that these two steps can be performed in parallel for each zone. At each ADMM iteration, the score of each trajectory is then computed from the pre-computed trajectory with respect to the current power targets $\bar{\textbf{u}}_{i}$.
III-B3 Maximum power constraint conversion